OpenStudy (anonymous):

trying to understand finding exact solutions on the interval 0, 2pi. when cos t= -.567? I know the answer is 2.17 but how do i find other possible exact answers?

OpenStudy (anonymous):

so is there a positive and negative answer in radians?

OpenStudy (anonymous):

but by taking the negative inverse of the cosine value (-.567) that's how I got the first result

OpenStudy (campbell_st):

well thats not quite right... as if you have you calculator in radian mode find cos(2.17) what happens...?

OpenStudy (anonymous):

i get -.5639

OpenStudy (campbell_st):

ok... so thats a 2nd quadrant angle... 2.17 radians... the 3rd quadrant where cos is also negative is \[\pi + 2.17 =\]

OpenStudy (anonymous):

well that equals 5.31?

OpenStudy (campbell_st):

so check by finding cos(5.31) what do you get...?

OpenStudy (anonymous):

.9957

OpenStudy (campbell_st):

ummm... somethings not right...

OpenStudy (campbell_st):

I just realised so find the angle where cos(t) = .567

OpenStudy (anonymous):

I'm not sure how to find that?

OpenStudy (campbell_st):

just inverse cos(0.567)

OpenStudy (anonymous):

i get .9679?

OpenStudy (campbell_st):

great so t = 0.9679 radians cos is negative in the 2nd and 3rd quadrants... so 2nd quadrant use \[\pi - t \] 3rd quadrant use \[\pi + t\] and you'll have your 2 angles...

OpenStudy (anonymous):

so I have 2.17 & 4.11 but shouldn't the 4.11 be a negative?

OpenStudy (campbell_st):

the 2nd quadrant angle will be 2.17.... as you were able to get with the calculator so to test the 3rd quadrant angle find cos(4.11) and it should be -0.567

OpenStudy (anonymous):

aaaahhh... ok. So how do you know when to add/subtract pi, or add/subtract 2pi?

OpenStudy (campbell_st):

yep... the method is 1 find the value of the positive so find inverse cos(0.567) that gives the angle know the quadrants and which ratios are positive and negative.. then know 2nd quadrant angles are \[\pi - \theta\] 3rd quadrant angles are \[\pi + \theta\] 4th quadrant angles are \[2\pi - \theta\] hope it helps

OpenStudy (anonymous):

oh. great. ty very much. I appreciate your help. It's much clearer now.